# Fundamental computational limits of weak learnability in high-dimensional multi-index models

### v2

@@ Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks. This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$. The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that it is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge. Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones. Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
@@ 1 Introduction
A fundamental property of neural networks is their ability to learn features from data and adapt to relevant structures in high-dimensional noisy data. However, our mathematical understanding of this mechanism remains limited. A popular model for studying this question is multi-index models. Multi-index functions are a class of statistical models encoding the inductive bias that the relevant directions for prediction depend only on a low-dimensional subspace of the covariates ${\bf x}\in\mathbb{R}^{d}$:
@@
where the mapping $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ can further be stochastic i.e. depend on independent noise $\xi\in\mathbb{R}^{p^{\prime}}$ for some finte $p^{\prime}$. They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2023).
@@
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020) in the same regime. Similar results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Damian et al. (2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
@@
  • We analyse the computational limits of weak learnability (as defined in Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and constant number of indices $p=\Theta(1)$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$. When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small initial overlap $\epsilon>0$ in $\mathcal{O}(\log\frac{1}{\epsilon})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition. Heuristically, we expect recovery along easy directions from random initialization with $\alpha>\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations. We further confirm this numerically through extensive simulations.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. It is conjectured that no efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
@@
2 Settings and definitions
As motivated in the introduction, our main focus in this work will be to study subspace identifiability in the class of Gaussian multi-index models.
@@ 2 Settings and definitions
where $g:\mathbb{R}^{p}\to\mathbb{R}$ denotes the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\xi}_{2}^{2}\right]<\infty$. [^3] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise. But for simplicity of exposition we focus on the $p^{\prime}=p$ case.
Given $n$ i.i.d. samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$ drawn as per Definition 1, we are interested in investigating the computational bottlenecks of estimating ${{\bm{{W}}}}^{\star}$ from the samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$. Note that reconstructing ${{\bm{{W}}}}^{\star}$ or a permutation of its rows is equivalent from the perspective of the likelihood Equation 2. Therefore, in this work, we will be interested in weak subspace learnability, which corresponds to obtaining an estimation of the subspace spanned by ${{\bm{{W}}}}^{\star}$ better than a random estimator. This can be defined in an invariant way as follows
@@
Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem. Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ (defined later in Equation 4). Therefore, it belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$. Intuitively, AMP estimates the mean ${\bm{\omega}}$ and variance ${\bm{V}}$ of ${{\bm{{W}}}}^{*}{\bm{x}}$ and the mean $\hat{{{\bm{{W}}}}}$ and covariance $\hat{{\bm{C}}}$ of ${{\bm{{W}}}}^{*}$ through a fixed point iteration described in Algorithm 1. The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020). This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
@@
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$. Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations.
Our results rely on the following assumption on the above denoiser, translating to an indirect assumption on $g,\xi$:
@@
${\bm{g}}_{\rm out}:\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\in\mathcal{C}^{2}$. Furthermore, the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},\xi)={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},\xi),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$. Here, $\xi\in\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\xi}^{2}\right]<\infty$ as per Definition 1.

### v3

@@ Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural nets. This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples $n\!=\!\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!\!>\!\!0$; (ii) if the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace where directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$, where $\alpha_{c}$ marks a computational phase transition. In a limited but interesting set of really hard directions –akin to the parity problem– $\alpha_{c}$ is found to diverge. Finally, (iii) we show that interactions between different directions can result in an intricate hierarchical learning phenomenon, where directions can be learned sequentially when coupled to easier ones. We discuss in detail the grand staircase picture associated to these functions (and contrast it with the original staircase one). Our theory builds on the optimality of approximate message-passing among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent, which we discuss in this context.
@@ 1 Introduction
A fundamental property of neural networks is their ability to learn features and adapt to relevant structures in high-dimensional noisy data. However, our mathematical understanding of this mechanism remains limited. A popular model for studying this question is multi-index models. Multi-index functions are a class of statistical models encoding the inductive bias that the relevant directions for prediction depend only on a low-dimensional subspace of the covariates ${\bf x}\in\mathbb{R}^{d}$:
@@
where the mapping $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ can further be stochastic i.e. depend on independent noise $\xi\in\mathbb{R}^{p^{\prime}}$ for some finte $p^{\prime}$. They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is used when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2023).
@@
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from the model in Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p\!=\!1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were also derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020). Algorithmic results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Here $\ell$ denotes the information exponent (Arous et al., 2021) of the target defined as the first non-zero integer $\ell\!>\!0$ in the Hermite expansion of $g$: $\ell\!=\!\min\{j\!\in\!\mathbb{N}:\langle g,\mathcal{H}_{j}\rangle_{\gamma}\neq 0\}.$
A related notion is that of staircase functions (Abbe et al., 2022; 2023), which characterizes the sample complexity for sequential learning of directions under the under the CSQ model or online SGD. For staircase functions, the CSQ sample complexity is governed by the “leap”, which informally equals the maximum jump in degree (in the Hermite expansion) conditioned on previously learned directions. For instance the target $g(z_{1}z_{2}z_{3})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}$ has leap $1$, since the term $z_{1}z_{2}$ is linear conditioned on $z_{1}$ while $z_{1}z_{2}z_{3}$ is linear conditioned on $z_{1},z_{2}$. Additionally Damian et al. (2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, and a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent. As we shall see, our results allows to reconcile this picture with an extended version of staircase functions that we refer to as grand staircase functions.
@@
  • We analyse the computational limits of weak learnability (Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\!\to\!\infty$ at fixed ratio $\alpha\!=\!\nicefrac{{n}}{{d}}$ and constant number of indices $p\!=\!\Theta(1)$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural nets trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples: trivial directions can be learned with a single AMP iteration for any $\alpha\!>\!0$. When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small side-information $\sqrt{\lambda}\!>\!0$ in $\mathcal{O}(\log\nicefrac{{1}}{{\lambda}})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition. Heuristically, we expect recovery along easy directions from random initialization with $\alpha\!>\!\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations and confirm this numerically through extensive simulations.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. It is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena, with the hierarchy referring to learning a sequence of growing subspaces. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We characterise this class of grand staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
Finally, the code to reproduce our plots, to run the AMP algorithm, and to deploy state evolution is available on GitHub https://github.com/SPOC-group/FundamentalLimitsMultiIndex
@@ 2 Settings and definitions
where $g:\mathbb{R}^{p}\to\mathbb{R}$ denotes the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\bm{\xi}}_{2}^{2}\right]<\infty$. [^4] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise. But for simplicity of exposition we focus on the $p^{\prime}=p$ case.
Given $n$ i.i.d. samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$ drawn as per Definition 1, we are interested in investigating the computational bottlenecks of estimating ${{\bm{{W}}}}^{\star}$ from the samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$. Note that reconstructing ${{\bm{{W}}}}^{\star}$ or a permutation of its rows is equivalent from the perspective of the likelihood Equation 2. Therefore, in this work, we will be interested in weak subspace learnability, which corresponds to obtaining an estimation of the subspace spanned by ${{\bm{{W}}}}^{\star}$ better than a random estimator. This can be defined in an invariant way:
@@
Our main tool for characterising the computational bottlenecks in the multi-index problem is an Approximate Message Passing (AMP, Algorithm 1) acting on ${\bf B}\in\mathbb{R}^{d\times p}$ and ${\bf\Omega}\in\mathbb{R}^{n\times p}$ tailored to our needs:
${\bm{\Omega}}^{t}$ $={\bm{X}}f_{t}({\bf B}^{t})-g_{t-1}({\bm{\Omega}}^{t-1},{\bf y}){\bm{V}}_{t}$ (4)
${\bf B}^{t+1}$ $={\bm{X}}^{T}g_{t}({\bm{\Omega}}^{t},{\bf y})+f_{t}({\bf B}^{t}){\bm{A}}_{t}$ (5)
where the denoisers $g_{t}(.,y_{i})$ and $f_{t}(.)$ are vector-valued mapping applied row-wise on ${\bf\omega}_{i}\in\mathbb{R}^{p}={\bf\Omega}_{.i}$ and ${\bf b}_{i}\in\mathbb{R}^{p}={\bf B}_{.i}$. Here $\hat{{\bm{{W}}}}^{t}\!\in\!\mathbb{R}^{p\times d}=f_{t}({\bf B}^{t})^{\top}$ are the estimator of the weights matrix ${{\bm{{W}}}}^{\star}$, $g_{t}({\bm{\Omega}}^{t})\in\mathbb{R}^{n\times p}$ are an estimate of the pre-activations, while ${\bm{A}}_{t}$ and ${\bm{V}}_{t}$ are the so-called Onsager terms (see Sec.A).
AMP only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and thus belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$. The key property of AMP is that for well-chosen $g_{t}$ and $f_{t}$ (using Bayesian denoisers) it is provably optimal within the class of first-order methods (Celentano et al., 2020). In our case, this require $g_{t}({\bm{\omega}},y)={\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ as in eq. (6), while $f_{t}$ is just a Gaussian Bayesian denoiser (see Sec.A). Learnability for the optimal AMP thus implies a computational lower bound in the class of first-order methods, including in particular popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, ${\bm{g}}_{\rm out}$ is given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$, which reads
@@
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$. Our results rely on the following assumption on the above denoiser, translating to weak assumptions on $g,\xi$:
@@
(a) ${\bm{g}}_{\rm out}\!:\!\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\!\in\!\mathcal{C}^{2}$. (b) the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},{\bm{\xi}})={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},{\bm{\xi}}),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$. Here, ${\bm{\xi}}\!\in\!\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\bm{\xi}}^{2}\right]\!<\!\infty$ as per Def. 1.